OpenStudy (anonymous):

How much energy is needed to raise the temperature of 40.0 g of argon from 25°C to 40°C? The specific heat capacity of argon is 0.520 J/(gK). A 20.8 J B 208 J C 416 J D 312 J

OpenStudy (sreekar369):

Q=mst(where t is change in temprature) Q=40*0.52*(40-25) Q=40*0.52*15 Q=312J(so its option d)

OpenStudy (sreekar369):

@Paule123